# VizAbility: Enhancing Chart Accessibility with LLM-based Conversational Interaction

### v2

@@ 6. Evaluation: User Study with Blind People
During the development process, we engaged with a blind participant who had prior experience using a screen reader on a daily basis. This participant, as a design partner, provided feedback at two intermediate stages of development. In addition to this intermediate prototype evaluation, we conducted a formal usability study with six additional blind/low-vision individuals. IRB approval was obtained before the commencement of the study.
@@ 6.4.2. Exploring the conversational module
Following this independent exploration of the conversational model, participants were made aware of the four distinct types of queries and were once again directed to input their own questions; however, this time around, they had to broadly adhere to one of the 4 query classifications. Users demonstrated a greater proficiency with the conversational module during this guided exploration, with P1 even chaining multiple individual queries to arrive at a broader understanding of the chart. By consecutively asking “What is the temperature for 2020?” and “What color is 2020?”, the participant was able to deduce that the color ‘orange-red’ indicates positive temperature values.
@@ 6.4.3. Integrating the two components
Participants were then introduced to navigation queries. We explained the purpose of these queries, emphasizing their role in wayfinding and orientation, and then allowed them to formulate their own navigation queries. All users concurred that these queries were essential for understanding the tree view (P6: ”Maybe not as much as for smaller datasets, but I definitely see their use for complex data like this”), a sentiment echoed in the usefulness ratings they assigned to the integrated system. While previous ratings for the individual components averaged around 3, after this introduction, participants consistently rated the complete system between 4 and 5, with 5 being extremely useful.
Most participants tended to input short and concise navigation queries. Rather than inputting “How do I get from my current location to the percentage vaccinated value for Guam”, one user (P5) opted for the much simpler “Take me to Guam”. Showcasing its conversational strengths, our model was able to precisely identify the starting as well as ending nodes from this colloquial text, yielding the instructions: “Press the right arrow key. Press the down arrow key. Press the down arrow key.”
@@ 6.6. Changes After User Feedback
The extensive participant feedback we collected is reflected in our design decisions (Section 3) and iterative changes to the system (Section 4). Here, we briefly describe how we incorporated the lessons learned from the user study.
@@
—Participant 4’s experience revealed a key issue in our contextual query handling: the LLM often misinterpreted user intent due to a lack of information about the chart. For instance, it misunderstood P4’s question about “temporal polarity” as a linguistic term. To address this issue, we enhanced the background information by incorporating the tree view text of the chart into the Web Browser Agent’s prompt. This change led to better responses, such as correctly interpreting the term in the global temperature dataset. This observation also prompted us to further investigate the ambiguity in user queries, leading to the refinement of questions prior to processing (see Section 4.5).
@@
—Participant 2’s preference for automatic traversal, prompted by the output of a navigation query, highlighted that manually pressing a series of keys can be cumbersome, especially for lengthy navigation paths. Originally, we aimed to ensure transparency and grant users control over the process. However, recognizing this issue, we decided to introduce an option for automatic traversal (see Section 4.4.4). In practice, as users become familiar with the logistics and behavior of the navigation query, they might find manual traversal burdensome. Nevertheless, we have retained this feature as an optional setting to cater to users with different levels of technical proficiency.
@@
—To address Participant 5’s concerns regarding the complexity of navigating the tree view, we introduced a conventional raw data table as an alternative (see 4.3). Though this addition may not significantly contribute to the novelty of our work, it underscores our dedication to creating an inclusive system.
@@
—The initial self-guided exploration of charts showed that most participants (P1, P2, P3, P5) struggled to ask questions beyond data-retrieval queries. Despite recognizing the value of visual, contextual, and navigation queries, as indicated by their high ratings, participants were unaware of these query types until they were explicitly explained by the moderator. This observation, along with P2’s suggestion for help documentation and preference for interactive guidance, led to the addition of query suggestions to the system (see Section 4.5 and Section 4.6).
@@ 7.1. Limitations and Opportunities
Despite our initial aim to offer concise and informative answers, P5’s recommendation for user-adjustable response verbosity underscored the importance of user agency over designer-imposed settings. Given that speech is processed serially, the text length read by screen readers becomes a pivotal design consideration. This concern has been reiterated in prior research (Zong et al., 2022; Ault et al., 2002; w3c, 2023; Jung et al., 2021). Similarly, offering users the capability to customize node descriptions in the tree view could prove advantageous.
@@
7.2. Need for Rigorous Benchmark Testing
The cornerstone of our project is the conversational module, designed to address the inherent limitations of keyboard navigation. While the existing dataset enabled a meaningful evaluation of response quality based on real-world queries, our study revealed the need for a more extensive benchmarking dataset.
@@ 7.2. Need for Rigorous Benchmark Testing
Our evaluation was constrained not only by the four chart types but also by the limited range of questions (e.g., three query types), preventing a full assessment of VizAbility’s capabilities. For example, although we noted some visual queries necessitating reasoning over data and contextual queries seeking chart information, our limited dataset was not sufficient to capture multi-category questions requiring advanced multi-hop reasoning (Chen et al., 2020). Furthermore, our study did not evaluate situational questions related to a user’s current interest point within the tree view, which our dataset lacks. Questions dependent on understanding previous conversational context were also not explored. Considering the generative abilities of LLMs, synthetically generating these types of questions using human-created examples and advanced prompt engineering might be a viable method.
Performing comprehensive comparative studies with varied metrics
In our study, we compared our system exclusively with another that also assumes the availability of chart specifications, emphasizing reasoning over image understanding. While recent vision-based question-answering systems like ChartQA (Hoque et al., 2022) are noteworthy, public chatbots like ChatGPT, Bing, and Bard have also started supporting image inputs. Although these systems are still in the early stages of understanding synthetic images, such as graphic designs and data visualizations, beyond natural scenes (Bylinskii et al., 2017), a comparison with VizAbility could be insightful. A balanced evaluation approach might involve using an independent image parser to feed data into VizAbility, thereby concentrating on reasoning capabilities. A crucial aspect of this benchmarking lies in defining evaluation criteria that extend beyond mere correctness. These should include measures of clarity, conciseness, and relevance to the query.
@@ 7.3. Integrating into Existing Visualization Tools
Since VizAbility operates under the assumption that a chart specification is available, it may not be directly applicable to charts currently found on the web. Instead, our vision is to integrate VizAbility within existing data visualization platforms. Prior research underscores that many data visualization practitioners base their choices on the accessibility features of these platforms (Joyner et al., 2022). Another study highlights the lack of accessible design support these tools offer (Kim et al., 2023a). Exploring the design space to determine how VizAbility can seamlessly fit into current data visualization workflows would be compelling. Additionally, considering the degree of customization for data visualization designers, such as setting default verbosity levels, warrants further investigation.
@@ 8. Conclusion
In this study, we introduced VizAbility, a tool that enhances structured chart navigation through conversational interactions. Our quantitative assessments demonstrate a notable advancement beyond existing systems, and our qualitative analyses underscore the importance of the system’s integrated approach and its dedication to transparency. As a direction for future research, we aim to develop an extensive benchmark dataset to foster continuous enhancements of VizAbility.

### v3

@@ 6. Evaluation: User Study with Blind People
During the development process, we engaged with a blind participant who had prior experience using a screen reader on a daily basis. This participant provided feedback at two intermediate stages of development in what was an iterative design process. In addition to this intermediate prototype evaluation, we conducted a preliminary usability study with six additional blind/low-vision individuals. IRB approval was obtained before the commencement of the study. The stimuli used in the study was an early version of VizAbility. While the major features remain unchanged, the improvements made after the study are described in Section 6.6.
@@ 6.4.2. Exploring the conversational module
Following this independent exploration of the conversational model, participants were made aware of the four distinct types of queries and were once again directed to input their own questions; however, this time around, they had to broadly adhere to one of the four query classifications. Users demonstrated a greater proficiency with the conversational module during this guided exploration, with P1 even chaining multiple individual queries to arrive at a broader understanding of the chart. By consecutively asking “What is the temperature for 2020?” and “What color is 2020?”, the participant was able to deduce that the color ‘red’ indicates positive temperature values.
@@ 6.4.3. Integrating the two components
Participants were then introduced to navigation queries. We explained the purpose of these queries, emphasizing their role in wayfinding and orientation, and then allowed them to formulate their own navigation queries. All users concurred that these queries were essential for understanding the tree view (P6: ”Maybe not as much as for smaller datasets, but I definitely see their use for complex data like this”), a sentiment echoed in the usefulness ratings they assigned to the integrated system. While previous Likert scale ratings for the individual components averaged around 3, after this introduction, participants consistently rated the complete system between 4 and 5, with 5 being extremely useful.
Most participants tended to input short and concise navigation queries. Rather than inputting “How do I get from my current location to the percentage vaccinated value for Guam”, one user (P5) opted for the much simpler “Take me to Guam”. Showcasing its conversational strengths, our model was able to precisely identify the starting as well as ending nodes from this colloquial text, yielding the instructions: “Press the right arrow key. Press the down arrow key two times.”
@@ 6.6. Changes After User Feedback
Here, we briefly describe how we incorporated the lessons learned from the user study.
@@
Participant 4’s experience revealed a limitation in our initial contextual query handling: VizAbility often misinterpreted user intent due to missing chart information. For instance, it treated “temporal polarity” as a linguistic term, not a data dimension depicted in the chart. We addressed this by enriching prompts with chart tree view text (Figure 3), which led to correct interpretation of the data (e.g., ”temporal polarity” in global temperature data). This observation also led to exploring query ambiguity in general, resulting in pre-processing question refinement (Section 4.5).
@@
Participant 2’s preference for automatic traversal suggested that manually pressing a series of keys can be cumbersome, especially for lengthy navigation paths. Originally, we aimed to ensure transparency and grant users control over the process. However, recognizing this issue, we decided to introduce an option for automatic traversal (see Section 4.4.4). While manual navigation might become tedious with familiarity, we kept it as a default option to accommodate varying technical proficiencies.
@@
To address Participant 5’s concerns regarding the complexity of navigating the tree view, we introduced a conventional raw data table as an alternative (see 4.3). Though this addition may not significantly contribute to the novelty of our work, it underscores our dedication to creating an inclusive system.
@@
The initial self-guided exploration of charts showed that most participants (P1, P2, P3, P5) struggled to ask questions beyond data-retrieval queries. Despite recognizing the value of visual, contextual, and navigation queries, participants were unaware of these query types until they were explicitly explained by the moderator. This observation, along with P2’s suggestion for help documentation and preference for interactive guidance, led to the addition of query suggestions to the system (see Section 4.5 and Section 4.6).
@@ 7.1. Limitations and Opportunities
Despite our initial aim to offer concise and informative answers, P5’s recommendation for user-adjustable response verbosity underscored the importance of user agency over designer-imposed settings. Given that speech is processed serially, the text length read by screen readers becomes a pivotal design consideration. This concern has been reiterated in prior research (Zong et al., 2022; Ault et al., 2002; w3c, 2023; Jung et al., 2022). Similarly, offering users the capability to customize node descriptions in the tree view could prove advantageous.
@@
7.2. Need for Rigorous and Inclusive Benchmark Testing
The cornerstone of our work is the conversational module, designed to address the challenges with keyboard navigation. While the existing dataset enabled a meaningful evaluation of response quality based on real-world queries, our study revealed the need for a more extensive and inclusive benchmarking dataset that incorporates viewpoints of blind and low-vision individuals (Gadiraju et al., 2023).
@@ 7.2. Need for Rigorous and Inclusive Benchmark Testing
Our evaluation was constrained not only by the four chart types but also by the limited range of questions (e.g., three query types), preventing a full assessment of VizAbility’s capabilities. For example, although we noted some visual queries necessitating reasoning over data and contextual queries seeking chart information, our limited dataset was not sufficient to capture multi-category questions requiring advanced multi-hop reasoning (Chen et al., 2020). Furthermore, our study did not evaluate situational questions related to a user’s current interest point within the tree view, which our dataset lacks. Questions dependent on understanding previous conversational context were also not explored. Considering the generative abilities of LLMs, synthetically generating these types of questions using human-created examples and advanced prompt engineering might be a viable method (Ko et al., 2023).
Incorporating vision and testing with varied metrics
Although GPT4V falls short in performance compared to VizAbility, it shows promising visual description capabilities. This advancement contrasts with a few years ago when interpreting synthetic images like graphic designs and data visualizations was inferior to natural scene images (Bylinskii et al., 2017). Similarly, recent image-based ChartQAs still depend on OCR to extract text from images and create data tables (Hoque et al., 2022; Kantharaj et al., 2022). Analyzing the capabilities of emerging vision-LLMs—from low-level analytic tasks like value look-ups and comparisons to higher-level cognitive operations like explaining chart patterns—will help find ways to integrate vision capability with the symbolic processing of VizAbility to achieve greater performance. Moreover, expanding beyond mere correctness to include other pertinent measures such as fluency, informativeness, and relevance to the query (Li et al., 2024) will be helpful for further improving the user experience of VizAbility.
@@ 7.3. Integrating into Existing Visualization Tools
Since VizAbility operates under the assumption that a chart specification is available, it may not be directly applicable to charts currently found on the web. Instead, our vision is to integrate VizAbility within existing data visualization platforms. Prior research underscores that many data visualization practitioners base their choices on the accessibility features of these platforms (Joyner et al., 2022). Another study highlights the lack of accessible design support these tools offer (Kim et al., 2023a). Exploring the design space to determine how VizAbility can seamlessly fit into current data visualization workflows would be compelling. Additionally, considering the degree of customization for data visualization designers, such as setting default verbosity levels and offering query guidance & suggestions, warrants further investigation.
@@ 8. Conclusion
In this study, we introduced VizAbility, a tool that enhances structured chart navigation through conversational interactions. Our quantitative assessments demonstrate a notable advancement beyond existing systems, and our qualitative analyses underscore the importance of the system’s integrated approach and its dedication to transparency. As a direction for future research, we aim to develop a more comprehensive and inclusive benchmark dataset to foster continuous enhancements of VizAbility, along with integrating vision capabilities to bridge the current performance gaps.
Acknowledgements.
[currently redacted for anonymous review]